OpenStudy (anonymous):

How do you find the polynomial function that has the given zeroes? 1 + radical 3, 1 - radical 3

jimthompson5910 (jim_thompson5910):

\[\Large x = 1+\sqrt{3} \ \text{or} \ x = 1-\sqrt{3}\] \[\Large x = 1\pm\sqrt{3}\] \[\Large x - 1 = \pm\sqrt{3}\] \[\Large (x - 1)^2 = 3\] \[\Large (x - 1)^2 - 3 = 0\] \[\Large x^2-2x+1- 3 = 0\] \[\Large x^2-2x-2= 0\] Polynomial function is \[\Large f(x) = x^2-2x-2\]